Operations Report – May 6th

Crew 235 Operations Report 6-MAY-2021

SOL: 4

Name of person filing report: Atila Meszaros

Non-nominal systems: ATV-2

Notes on non-nominal systems: ATV-2 gas pedal system is broken. Once you press it, it would get stuck, instead of retracting, and it could lead to an accident. There’s also a shift problem, but we suspect is related to the gas pedal system interfering with the shift change. As we leave on Saturday, the ATVs will have to be tuned up before next season.

Generator: Working nominally.

Hours run: 0 hours

From what time last night: the generator didn’t run

To what time this morning: the generator didn’t run

List any additional daytime hours when the generator was run: N/A

Solar— SOC 92% at 20:09 (Before generator is run at night)

Notes on power system: Nothing to report

Diesel Reading – 47%

Station Propane Reading – 80%

Outpost Propane Reading – 1. 70% 2. 80%

Water (loft tank): 25 gallons

Water Meter: 151161.3 units

Water (static tank): 250 gallons

Static to Loft Pump used – yes

Water in GreenHab: 0 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: No

Sojourner rover used: No, in town

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Spirit rover used: No, in town

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Opportunity rover used: No, in town

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Curiosity rover used: No, in town

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Perseverance rover used: Yes

Hours: 211.9

Beginning charge: 100%

Ending charge: 31%

Currently charging: Yes

Notes on rovers: None

ATV’s Used: Yes

Reason for use: EVA#4 and EVA#5

Oil Added? No

ATV Fuel Used: N/A

# Hours the ATVs were used today: 2 hours

Notes on ATVs: ATV 2 gs pedal out of service

HabCar used and why, where?: Yes, to go to town for supplies.

CrewCar used and why, where?: Nothing to report

General notes and comments: Nothing to report

Summary of internet: Working nominally.

Summary of suits and radios: All radios nominal, all functional suits currently charging.

Summary of Hab operations: Mouse found in the morning and released during EVA #5.

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: General cleaning and bacterial disposal.

Summary of RAM operations: Nothing to report.

Summary of any observatory issues: Battery replaced and now rotation system is working.

Summary of health and safety issues: Nothing to report

Questions, concerns, and requests to Mission Support: Thank you for being with us tonight, Bernard!